OpenStudy (anonymous):

I am practicing another integral. (check, but don't figure out what I should see for me)

OpenStudy (anonymous):

\[\large \int\limits_{ }^{ } \frac{x-1}{x^2+2}~dx\]

OpenStudy (anonymous):

wait, can I get a hint, but just a hint...

OpenStudy (jtvatsim):

Hint: u-substitution

OpenStudy (anonymous):

I was thinking of it, but which sub for what, for x-1?

OpenStudy (anonymous):

well, lets go that way.

OpenStudy (anonymous):

\[\large \int\limits_{ }^{ } \frac{u}{u^2-2x+3}~du\]

OpenStudy (jtvatsim):

Do you like what you see?

OpenStudy (anonymous):

well, not really, but seems a bit better than before.

OpenStudy (anonymous):

hold on, I can factor

OpenStudy (anonymous):

No I can't

OpenStudy (anonymous):

ohh, SHhhh the x, it is u

ganeshie8 (ganeshie8):

Hint : split the given integral into two integrals, one can be worked by simple u substitution and the other has an obvious trig substitution

OpenStudy (anonymous):

\[\large \int\limits_{ }^{ } \frac{x}{x^2+2}~dx-\int\limits_{ }^{ } \frac{1}{x^2+2}dx\]

OpenStudy (anonymous):

yes, the first one: u=x^2, (1/2)du=x dx

OpenStudy (anonymous):

Ohh, u=x62+2

OpenStudy (anonymous):

Yes, i thought of that

ganeshie8 (ganeshie8):

yes that substitution gives u a simpler integrand you need to look ahead of what you get by ur substition integration requires guessing

OpenStudy (anonymous):

\[\frac{1}{2}\large \int\limits_{ }^{ } \frac{1}{u}~du-\int\limits_{ }^{ } \frac{1}{x^2+2}dx\]\[\frac{1}{2}\ln \left| x^2+2 \right|-\int\limits_{ }^{ } \frac{1}{x^2+2}dx\]

ganeshie8 (ganeshie8):

looks good! for the other integral you may use table of integrals or try below trig substitution : \[x = \sqrt{2}\tan (\theta)\] and recall the trig identity \[\tan^2\theta + 1 = \sec^2\theta \]

OpenStudy (anonymous):

and 1/sqrt{2} d(theta) = dx ?

OpenStudy (anonymous):

\[\frac{1}{2}\ln \left| x^2+2 \right|-\frac{1}{2\sqrt{2}}\int\limits_{ }^{ } \frac{1}{ \tan \theta+1}d \theta\]

OpenStudy (anonymous):

I took out the 1/2 in the inegral

OpenStudy (anonymous):

wait, it is \[\frac{1}{2}\ln \left| x^2+2 \right|-\frac{1}{2\sqrt{2}}\int\limits_{ }^{ } \frac{1}{ \tan^2\theta+1}d \theta\]

ganeshie8 (ganeshie8):

Hmm, I suggest we use table of integrals..

OpenStudy (anonymous):

\[\frac{1}{2}\ln \left| x^2+2 \right|-\frac{1}{2\sqrt{2}}\tan^{-1}(\tan \theta +1)+C\]

OpenStudy (anonymous):

and then I do substitution

OpenStudy (anonymous):

or something wrong?

ganeshie8 (ganeshie8):

\[\large \int \dfrac{1}{x^2+a^2}dx = \frac{1}{a}\tan^{-1}\left(\frac{x}{a}\right)\]

OpenStudy (anonymous):

omg, it's 10:30 I haven't even noticed..... lol I got to go. I am really sorry

ganeshie8 (ganeshie8):

Hey we are done : \[\int\limits \frac{1}{x^2+2}dx = \frac{1}{\sqrt{2}} \tan^{-1}\left(\frac{x}{\sqrt2}\right)\]

ganeshie8 (ganeshie8):

we could have worked it easily using the earlier trig substitution but above formula is also good to memorize ^